The following consecutive readings were taken with a level and a 4-metre leveling staff on a continuously sloping ground at common intervals of 30 m:

0.855 (on A), 1.55 m, 2,335, 3.825, 0.455, 1.380, 2.055, 2.855, 3.455, 0.585, 1.015, 1.850, 2.755, 3.845 (on B).

 

The RL of A was 380.500. Make entries in a level book and apply the usual checks. Determine the gradient of AB.



1
Expert's answer
2022-02-15T14:52:01-0500



The given observed reading are gradually increasing initially but that they suddenly decrease after the 5th and 10th readings. This indicates that the instrument has shifted after 5th and 10th readings.

"\sum"BS- "\sum"FS= 1.895-11.125= -9.230


"\sum"Rise-"\sum"Fall=0-9.230=-9.230


LastRL-1st RL= 371.270=380.50=-9.230


Falling Gradient of AB="\frac{difference of level}{Horizontal distance}"


Falling Gradient of AB= "\frac{9.230}{360}" ="\frac{1}{39}" =1 in 39